# Securing the AI Supply Chain: What Can We Learn From Developer-Reported Security Issues and Solutions of AI Projects?

### v1

@@ 4.4. AI security data analysis
To perform the analysis for our RQs, we utilized a comprehensive dataset curated from the manually labeled dataset in Section 4.2 and additional security discussions without security keywords predicted using our best-performing classifier from Section 4.3. For RQ1, we examined AI security reporting practices by analyzing their temporal trends, references to official CVE-IDs (cve2024), and distribution across AI project domains and model origin types. These aspects can inform the evolution and the formal reporting process of security issues in the AI supply chain, how they correlate with real-world AI applications and the key practice of model reuse (qin2024adversarial; jiang2023empirical; taraghi2024deep). To analyze the reporting distribution, we obtained the project domains using the pipeline_tag attribute in the model cards (mitchell2019model), assigning a single task (e.g., text-generation) to each project. These tasks were then grouped under the domains listed on the HF Hub API (huggingface_api). For model origin types, we categorized models as either foundational (base) or derived (downstream) using the base_model attribute (mitchell2019model) and the public count of downstream usages, e.g., finetunes (huggingface_api).
@@
To address RQ2 and RQ3, we employed thematic analysis (cruzes2011recommended) to analyze the identified security discussions. This is similar to previous works  (gao2024documenting; gao2025ai) that studied qualitative attributes of Software Engineering and AI artifacts. From the inference results and our manually labeled data, we drew a sample size of 368 from HF and 385 from GH. These sample sizes ensured statistical representativeness, providing a 95% confidence level for estimating the proportion of security categories within each source with a 5% margin of error (robb1963w), i.e., the proportions observed in our sample closely approximate those in the entire population. During this sampling process, we also kept track of false positive records from our classifier to analyze classifier errors. We followed the below three steps to extract data for both research questions simultaneously, where for every discussion, we analyzed both the reported issues and the solutions. Following the thematic analysis guidelines (cruzes2011recommended; braun2006using), this analysis aims to identify semantic themes, focusing on the explicit content of developers’ discussions. We adopted a realist epistemological orientation, treating the data as a direct reflection of the real-world security challenges that developers face and ensuring our resulting taxonomy is empirically grounded in developer-reported realities.
@@
Data familiarization. The first author collected and examined the documents, then extracted the key points related to security issues and proposed solutions. During the process, CVE/CWE records were also extracted from NVD (nvd2025datafeeds) if the vulnerability was explicitly mentioned in the issue. These documents were captured in spreadsheets and shared with the second and third authors for revision. Disagreements were resolved by discussions and meetings.
@@
Codes generation and validation. To ensure analytical rigor, we employed an iterative peer debriefing process (lincoln1985naturalistic; creswell2016qualitative) for this step. Peer debriefing is a collaborative approach where the code book is reviewed by an external party who was not involved in the coding process. This method is well-suited for developing and refining a code book through discussion and consensus, which is distinct from other methods that rely on calculating inter-rater reliability against a predefined code book. Consequently, formal metrics (e.g., Cohen’s Kappa) were not calculated, as our goal was collaborative refinement rather than measuring independent agreement (creswell2016qualitative). The process started with the first author generating the initial codes by assigning multiple descriptive labels for each sample. A set of labels was generated to capture the security issue (RQ2), while a distinctive second set of labels captured the solution or mitigation strategy (RQ3). A code book was then developed to document the codes, representative samples, and authors’ notes for consistent tracking and references. The second author, uninvolved in the coding, acted as the peer debriefer and reviewed this code book, providing feedback on the clarity, distinctiveness, and representativeness of the codes. Regular discussions were held among all three authors to discuss the feedback, resolve disagreements, and iteratively refine the code book by merging, splitting, or renaming codes. The process was repeated multiple times until consensus was reached.
@@
Theme synthesis and finalization. Following the final code book development, the first author grouped the validated codes into sub-themes and themes that reflect their nature, following thematic analysis standard procedures (braun2006using; cruzes2011recommended). This synthesis process was also conducted in parallel, with one set of codes organized into the issue taxonomy (RQ2) and the other into the solution taxonomy (RQ3). Spreadsheets were used to organize the data, key points, codes, and themes to facilitate discussions. The sub-themes and themes were then reviewed and confirmed by all authors in multiple rounds of discussions, and the process ended when mutual agreement was reached. An illustrative example for the coding process from raw data to issue themes (RQ2) is provided in Figure 2, and the same process was also applied for solution themes (RQ3). The entire process from data familiarization to theme synthesis took approximately a total of 304 man-hours.
@@ 5.1. RQ1: What is the extent of security reporting in AI-enabled projects?
Error Analysis. Our test set ensured the model’s actual performance as it was randomly sampled from the entire corpus and reflected real-world scenarios. To further enhance reliability, we manually validated the classifier’s outputs during the random sampling process of our thematic analysis (Section 4.4). We identified 116 false positives (13.3%) and revealed two primary error sources: (1) security-related keywords usage in non-security contexts (e.g., TheBloke/dolphin-2.5-mixtral-8x7b-GPTQ-1 contains “Malware api call sequence embedding” when referring to malware logs extraction), and (2) security-adjacent discussions not describing actual vulnerabilities (e.g., cpacker/MemGPT-2037 discussed implementing authorization without raising security concerns). All false positives are available in our replication package (replication_package). While our classifier was effective, these limitations highlight the challenge of distinguishing security-specific discussions. Consequently, RQ1’s results are unlikely to be affected given the margin of errors and the false positive rate. Findings of RQ2 and RQ3 were derived from the analysis after false positives were removed, thus they are not impacted.
@@
The temporal distribution of the curated security reports illustrates an exponential increase over the years in AI security discourse (see Figure 3). The number of GH security reports more than doubled from 2022 to 2023, and reached a new peak of 83,561 in 2024. HF Discussion, launched late 2022 (hf2022disc), shows a similar explosive growth, increasing nearly tenfold from 487 reports in 2022 to 4,773 in 2024. This rapid growth in community discussions stands in stark contrast to the relatively small number of formally reported vulnerabilities. Only 320 (¡ 0.1%) discussions were linked to CVE-IDs, which are typically required for formal vulnerability reporting. This highlights the value of platforms like HF and GH in uncovering a broader range of real-world AI security issues that go beyond what is captured through traditional CVE-based tracking. To further characterize the AI-enabled projects where these security issues commonly arise, we analyzed the distribution of security reporting in terms of project domains and model origin types.
@@
Distribution of model origin types. Figure 5 reveals that base models are associated with a higher average number of security reports (2.55 per model) compared to downstream models (1.69 per model). These numbers suggest that foundation models attracted more security attention, likely due to their wider adoption and foundational role in the ecosystem (taraghi2024deep; qin2024adversarial). For example, the base google/gemma-2b, widely adopted with 23,536 usages, contained 80 discussions in total, while its fine-tuned google/gemma-2b-aps-it variant only contained three discussions on HF. However, downstream models still exhibit a substantial number of security issues, possibly arising from task-specific adaptations or integrations.
@@ 5.2. RQ2: What are common AI security issues?
As described in Section 4.4, we synthesized 32 issue codes in four main themes: (1) System and Software, (2) External Tools and Ecosystem, (3) Model, and (4) Data. The first two cover vulnerabilities in code and dependencies, echoing traditional software security but in the AI context; whereas, the Model and Data themes capture unique AI-specific surfaces, highlighting novel challenges in securing the AI supply chain. Table 3 presents our taxonomy with the total “Frequency” (the count of all code occurrences for each sub-theme) and the “Coverage” (the percentage of unique documents in which the sub-theme appeared at least once). Below are the codes, sub-themes, and themes in the order of their frequency, along with their description. Note that we only report the key issue codes due to the space limit; the full taxonomy can be found at (tax_viz).
@@
Since AI-enabled projects contain software components, they are also susceptible to vulnerabilities found in traditional software systems (papernot2018sok; chen2024security). System Runtime Compatibility Issues (325). This sub-theme documents how functional conflicts, such as file, framework, or dependency mismatches, can introduce security risks by causing crashes or instability. Being the most reported sub-theme, incompatibility is illustrated as not only a functional fault but a significant driver of security issues. For example, google/long-t5-local-large-1 reported a runtime crash due to an unrecognized file type, while comfyanonymous/ComfyUI-5334 documented process failures caused by an unsupported data type in a framework. In other cases, incompatibilities with the host environment led to severe instability, including “kernel failure” (oobabooga/text-generation-webui-2927).
Code Execution Security Issues (229). This sub-theme captures reports of established code execution vulnerabilities. Developers have frequently discussed vulnerabilities such as path traversal, XSS, and open redirects, and often linked them to formal CVE reports. For instance, OpenCodeInterpreter/OpenCodeInterpreter-30 identified an Open Redirect vulnerability with CVE-2024-4940. A particularly prominent issue, with 85 instances across both platforms, is unsafe file deserialization, which directly leads to arbitrary code execution. This finding is consistent with existing studies (casey2024large; soodmalicious; zhao2024models) and the official documentation (pythonpickle). For example, in Alpha-VLLM/Lumina-Next-T2I-1, developers repeatedly warned that using Python’s pickle format is “inherently insecure” and is a known vector for malware.
Deployment and Infrastructure Security Issues (182). This sub-theme includes the failure points during the deployment. Insecure default configurations and user misconfigurations are most common, where either the default setting is not secure, or users apply settings without understanding the security implications. For instance, in huggingface/transformers-22944, a user criticized the auto-download feature as a “security hole” that could potentially download trojan virus, or in danswer-ai/danswer-2444 where the wildcard CORS setting could lead to information theft. Another important issue, though less frequently reported, is unaware compromised deployments. comfyanonymous/ComfyUI-5165 noted that users blindly using the “-listen 0.0.0.0” command had unknowingly exposed “more than 1,000 instances to the public internet”, making them vulnerable to complete takeover.
@@
Model Input and Prompt Security Issues (63). This sub-theme relates to how models process user-provided data. The most frequently discussed issue is model input manipulations and evasions. This issue includes patterns of prompt injection, where crafted input manipulate the model into unintended actions. Users in huggingface/tokenizers-1458 reported that prompt templates can be “exploited by injecting control tokens disguised within Jinja templates to manipulate the AI assistant’s behavior, such as redefining system prompts.” Beyond text, this code also includes adversarial manipulation of non-text inputs, with a prime example of the “image-scaling attacks” detailed in tensorflow/tensorflow-62334 where an input image could be manipulated to deceive the model. Another issue is insecure prompt template design, facilitating injection attacks. In nvidia/Llama3-ChatQA-1.5-70B-1, developers reported templates that use common strings (‘User:’) as delimiters instead of unique tokens, allowing an attacker to craft inputs that the model misinterprets as a system instruction. Another example is turboderp/Mistral-Nemo-Instruct-12B-exl2-1, where a “broken tokenizer” and “inconsistent prompt formatting” led to “persona shifts” and degraded output. These inconsistencies can cause reliability and security risks by making the model more susceptible to prompt injection or generating misleading or unauthorized responses (liu2024exploring).
@@
Another unique aspect of AI-enabled systems is their heavy reliance on a large volume of data that carries their own vulnerabilities (whang2023data). Our sub-themes capture how data can be exposed and how its (low) quality can affect the integrity of a model or a system.
@@ 5.3. RQ3: What are the solutions to these issues?
This section presents solutions to the identified security issues, ranging from concrete implementations to developer suggestions and best practices. We have synthesized 24 solution codes in a taxonomy parallel to RQ2 to facilitate a high-level comparison. Our analysis has revealed that solutions can be cross-cutting, addressing multiple issues across themes. We first describe the synthesized sub-themes as reported in Table 3, followed by the analysis of their relationships. Similar to RQ2, we only report the key solution codes due to the space limit; the full taxonomy is available at (tax_viz).
@@ 5.3.1. System and Software Security Solutions (369)
Secure Coding and Execution (153). These solutions focus on hardening the application source code, mainly to address Code Execution Issues. These include both reactive and proactive input validation, such as replacing insecure functions to prevent path traversal in microsoft/Olive-83 or proposing new utilities to validate URLs and prevent SSRF (gradio-app/gradio-8301). A critical solution has been the promotion of secure deserialization practices to mitigate risks from vulnerable formats. For instance, huggingface/transformers-27776 proposed gating pickle usage behind an explicit environment variable and Alpha-VLLM/Lumina-Next-T2I-1 advised the adoption of the safetensor (safetensors) format. More complex solutions to enable safer execution or implementations like sandboxing (deepset-ai/haystack-8691), omitting usage of system-level functions that can execute code (e.g., Python’s posix.system in mkiani/gpt2-system-1), or using time-constant methods to mitigate the potential side-channel attack on password (cpacker/MemGPT-1230) were also discussed. Yet, their adoptions were limited due to the high complexity.

### v2

@@ 4.4. AI security data analysis
To perform the analysis for our RQs, we utilized a comprehensive dataset curated from the manually labeled dataset in Section 4.2 and additional security discussions without security keywords predicted using our best-performing classifier from Section 4.3. For RQ1, we examined AI security reporting practices by analyzing their temporal trends, references to official CVE-IDs (of Standards and (NIST), 2024), and distribution across AI project domains and model origin types. These aspects can inform the evolution and the formal reporting process of security issues in the AI supply chain, how they correlate with real-world AI applications and the key practice of model reuse (Qin et al., 2024; Jiang et al., 2023; Taraghi et al., 2024). To analyze the reporting distribution, we obtained the project domains using the pipeline_tag attribute in the model cards (Mitchell et al., 2019), assigning a single task (e.g., text-generation) to each project. These tasks were then grouped under the domains listed on the HF Hub API (Face, 2024). For model origin types, we categorized models as either foundational (base) or derived (downstream) using the base_model attribute (Mitchell et al., 2019) and the public count of downstream usages, e.g., finetunes (Face, 2024).
@@
To address RQ2 and RQ3, we employed thematic analysis (Cruzes and Dyba, 2011) to analyze the identified security discussions. This is similar to previous works  (Gao et al., 2024, 2025) that studied qualitative attributes of Software Engineering and AI artifacts. From the inference results and our manually labeled data, we drew a sample size of 368 from HF and 385 from GH. These sample sizes ensured statistical representativeness, providing a 95% confidence level for estimating the proportion of security categories within each source with a 5% margin of error (Cochran, 2007), i.e., the proportions observed in our sample closely approximate those in the entire population. During this sampling process, we also kept track of false positive records from our classifier to analyze classifier errors. We followed the below three steps to extract data for both research questions simultaneously, where for every discussion, we analyzed both the reported issues and the solutions. Following the thematic analysis guidelines (Cruzes and Dyba, 2011; Braun and Clarke, 2006), this analysis aims to identify semantic themes, focusing on the explicit content of developers’ discussions. We adopted a realist epistemological orientation, treating the data as a direct reflection of the real-world security challenges that developers face and ensuring our resulting taxonomy is empirically grounded in developer-reported realities.
@@
Data familiarization. The first author collected and examined the documents, then extracted the key points related to security issues and proposed solutions. During the process, CVE/CWE records were also extracted from NVD (of Standards and (NIST), 2025) if the vulnerability was explicitly mentioned in the issue. These documents were captured in spreadsheets and shared with the second and third authors for revision. Disagreements were resolved by discussions and meetings.
@@
Codes generation and validation. To ensure analytical rigor, we employed an iterative peer debriefing process (Lincoln, 1985; Creswell and Poth, 2016) for this step. Peer debriefing is a collaborative approach where the code book is reviewed by an external party who was not involved in the coding process. This method is well-suited for developing and refining a code book through discussion and consensus, which is distinct from other methods that rely on calculating inter-rater reliability against a predefined code book. Consequently, formal metrics (e.g., Cohen’s Kappa) were not calculated, as our goal was collaborative refinement rather than measuring independent agreement (Creswell and Poth, 2016). The process started with the first author generating the initial codes by assigning multiple descriptive labels for each sample. A set of labels was generated to capture the security issue (RQ2), while a distinctive second set of labels captured the solution or mitigation strategy (RQ3). A code book was then developed to document the codes, representative samples, and authors’ notes for consistent tracking and references. The second author, uninvolved in the coding, acted as the peer debriefer and reviewed this code book, providing feedback on the clarity, distinctiveness, and representativeness of the codes. Regular discussions were held among all three authors to discuss the feedback, resolve disagreements, and iteratively refine the code book by merging, splitting, or renaming codes. The process was repeated multiple times until consensus was reached.
@@
Theme synthesis and finalization. Following the final code book development, the first author grouped the validated codes into sub-themes and themes that reflect their nature, following thematic analysis standard procedures (Braun and Clarke, 2006; Cruzes and Dyba, 2011). This synthesis process was also conducted in parallel, with one set of codes organized into the issue taxonomy (RQ2) and the other into the solution taxonomy (RQ3). Spreadsheets were used to organize the data, key points, codes, and themes to facilitate discussions. The sub-themes and themes were then reviewed and confirmed by all authors in multiple rounds of discussions, and the process ended when mutual agreement was reached. An illustrative example for the coding process from raw data to issue themes (RQ2) is provided in Figure 2, and the same process was also applied for solution themes (RQ3). The entire process from data familiarization to theme synthesis took approximately a total of 304 man-hours.
@@ 5.1. RQ1: What is the extent of security reporting in AI-enabled projects?
Error Analysis. Our test set ensured the model’s actual performance as it was randomly sampled from the entire corpus and reflected real-world scenarios. To further enhance reliability, we manually validated the classifier’s outputs during the random sampling process of our thematic analysis (Section 4.4). We identified 116 false positives (13.3%) and revealed two primary error sources: (1) security-related keywords usage in non-security contexts (e.g., TheBloke/dolphin-2.5-mixtral-8x7b-GPTQ-1 contains “Malware api call sequence embedding” when referring to malware logs extraction), and (2) security-adjacent discussions not describing actual vulnerabilities (e.g., cpacker/MemGPT-2037 discussed implementing authorization without raising security concerns). All false positives are available in our replication package (Nguyen, 2025b). While our classifier was effective, these limitations highlight the challenge of distinguishing security-specific discussions. Consequently, RQ1’s results are unlikely to be affected given the margin of errors and the false positive rate. Findings of RQ2 and RQ3 were derived from the analysis after false positives were removed, thus they are not impacted.
@@
The temporal distribution of the curated security reports illustrates an exponential increase over the years in AI security discourse (see Figure 3). The number of GH security reports more than doubled from 2022 to 2023, and reached a new peak of 83,561 in 2024. HF Discussion, launched late 2022 (Hugging Face, 2022), shows a similar explosive growth, increasing nearly tenfold from 487 reports in 2022 to 4,773 in 2024. This rapid growth in community discussions stands in stark contrast to the relatively small number of formally reported vulnerabilities. Only 320 (¡ 0.1%) discussions were linked to CVE-IDs, which are typically required for formal vulnerability reporting. This highlights the value of platforms like HF and GH in uncovering a broader range of real-world AI security issues that go beyond what is captured through traditional CVE-based tracking. To further characterize the AI-enabled projects where these security issues commonly arise, we analyzed the distribution of security reporting in terms of project domains and model origin types.
@@
Distribution of model origin types. Figure 5 reveals that base models are associated with a higher average number of security reports (2.55 per model) compared to downstream models (1.69 per model). These numbers suggest that foundation models attracted more security attention, likely due to their wider adoption and foundational role in the ecosystem (Taraghi et al., 2024; Qin et al., 2024). For example, the base google/gemma-2b, widely adopted with 23,536 usages, contained 80 discussions in total, while its fine-tuned google/gemma-2b-aps-it variant only contained three discussions on HF. However, downstream models still exhibit a substantial number of security issues, possibly arising from task-specific adaptations or integrations.
@@ 5.2. RQ2: What are common AI security issues?
As described in Section 4.4, we synthesized 32 issue codes in four main themes: (1) System and Software, (2) External Tools and Ecosystem, (3) Model, and (4) Data. The first two cover vulnerabilities in code and dependencies, echoing traditional software security but in the AI context; whereas, the Model and Data themes capture unique AI-specific surfaces, highlighting novel challenges in securing the AI supply chain. Table 3 presents our taxonomy with the total “Frequency” (the count of all code occurrences for each sub-theme) and the “Coverage” (the percentage of unique documents in which the sub-theme appeared at least once). Below are the codes, sub-themes, and themes in the order of their frequency, along with their description. Note that we only report the key issue codes due to the space limit; the full taxonomy can be found at (Nguyen, 2025a).
@@
Since AI-enabled projects contain software components, they are also susceptible to vulnerabilities found in traditional software systems (Papernot et al., 2018; Chen and Babar, 2024). System Runtime Compatibility Issues (325). This sub-theme documents how functional conflicts, such as file, framework, or dependency mismatches, can introduce security risks by causing crashes or instability. Being the most reported sub-theme, incompatibility is illustrated as not only a functional fault but a significant driver of security issues. For example, google/long-t5-local-large-1 reported a runtime crash due to an unrecognized file type, while comfyanonymous/ComfyUI-5334 documented process failures caused by an unsupported data type in a framework. In other cases, incompatibilities with the host environment led to severe instability, including “kernel failure” (oobabooga/text-generation-webui-2927).
Code Execution Security Issues (229). This sub-theme captures reports of established code execution vulnerabilities. Developers have frequently discussed vulnerabilities such as path traversal, XSS, and open redirects, and often linked them to formal CVE reports. For instance, OpenCodeInterpreter/OpenCodeInterpreter-30 identified an Open Redirect vulnerability with CVE-2024-4940. A particularly prominent issue, with 85 instances across both platforms, is unsafe file deserialization, which directly leads to arbitrary code execution. This finding is consistent with existing studies (Casey et al., 2024; Sood and Zeadally, 2025; Zhao et al., 2024a) and the official documentation (Foundation, ). For example, in Alpha-VLLM/Lumina-Next-T2I-1, developers repeatedly warned that using Python’s pickle format is “inherently insecure” and is a known vector for malware.
Deployment and Infrastructure Security Issues (182). This sub-theme includes the failures during deployment. Insecure default configurations and user misconfigurations (Ye et al., 2025) are most common, where either the default setting is not secure, or users apply settings without understanding the security implications. For instance, in huggingface/transformers-22944, a user criticized the auto-download feature as a “security hole” that could download trojan virus, or in danswer-ai/danswer-2444 where the wildcard CORS setting could lead to information theft. Another important issue, though less frequently reported, is unaware compromised deployments. comfyanonymous/ComfyUI-5165 noted that users blindly using the “-listen 0.0.0.0” command had unknowingly exposed “more than 1,000 instances to the public internet”, making them vulnerable to complete takeover.
@@
Model Input and Prompt Security Issues (63). This sub-theme relates to how models process user-provided data. The most frequently discussed issue is model input manipulations and evasions. This issue includes patterns of prompt injection, where crafted input manipulate the model into unintended actions. Users in huggingface/tokenizers-1458 reported that prompt templates can be “exploited by injecting control tokens disguised within Jinja templates to manipulate the AI assistant’s behavior, such as redefining system prompts.” Beyond text, this code also includes adversarial manipulation of non-text inputs, with a prime example of the “image-scaling attacks” detailed in tensorflow/tensorflow-62334 where an input image could be manipulated to deceive the model. Another issue is insecure prompt template design, facilitating injection attacks. In nvidia/Llama3-ChatQA-1.5-70B-1, developers reported templates that use common strings (‘User:’) as delimiters instead of unique tokens, allowing an attacker to craft inputs that the model misinterprets as a system instruction. Another example is turboderp/Mistral-Nemo-Instruct-12B-exl2-1, where a “broken tokenizer” and “inconsistent prompt formatting” led to “persona shifts” and degraded output. These inconsistencies can cause reliability and security risks by making the model more susceptible to prompt injection or generating misleading or unauthorized responses (Liu and Hu, 2024).
@@
Another unique aspect of AI-enabled systems is their heavy reliance on a large volume of data that carries their own vulnerabilities (Whang et al., 2023). Our sub-themes capture how data can be exposed and how its (low) quality can affect the integrity of a model or a system.
@@ 5.3. RQ3: What are the solutions to these issues?
This section presents solutions to the identified security issues, ranging from concrete implementations to developer suggestions and best practices. We have synthesized 24 solution codes in a taxonomy parallel to RQ2 to facilitate a high-level comparison. Our analysis has revealed that solutions can be cross-cutting, addressing multiple issues across themes. We first describe the synthesized sub-themes as reported in Table 3, followed by the analysis of their relationships. Similar to RQ2, we only report the key solution codes due to the space limit; the full taxonomy is available at (Nguyen, 2025a).
@@ 5.3.1. System and Software Security Solutions (369)
Secure Coding and Execution (153). These solutions focus on hardening the application source code, mainly to address Code Execution Issues. These include both reactive and proactive input validation, such as replacing insecure functions to prevent path traversal in microsoft/Olive-83 or proposing new utilities to validate URLs and prevent SSRF (gradio-app/gradio-8301). A critical solution has been the promotion of secure deserialization practices to mitigate risks from vulnerable formats. For instance, huggingface/transformers-27776 proposed gating pickle usage behind an explicit environment variable and Alpha-VLLM/Lumina-Next-T2I-1 advised the adoption of the safetensor (Face, 2023) format. More complex solutions to enable safer execution or implementations like sandboxing (deepset-ai/haystack-8691), omitting usage of system-level functions that can execute code (e.g., Python’s posix.system in mkiani/gpt2-system-1), or using time-constant methods to mitigate the potential side-channel attack on password (cpacker/MemGPT-1230) were also discussed. Yet, their adoptions were limited due to the high complexity.